OpenStudy (anonymous):

Find the length of a if the measure of angle A equals 50 degrees, b = 20, and c = 25. Show work. Round answer to the nearest hundredth.

OpenStudy (jdoe0001):

well, you have angle "A", and it's facing off side would be "a" you also are given sides "b" and "c" "b" and "c" most likely converge in the angle that IS NOT "B" or "C", meaning, "A" so you have and angle and its 2 adjacents sides

OpenStudy (jdoe0001):

so :), what do you think?

OpenStudy (anonymous):

Okay so, you use the law of cosine

OpenStudy (jdoe0001):

yeah

OpenStudy (jdoe0001):

pretty straightforward, just plug in the numbers

OpenStudy (anonymous):

\[50^{2} \left( 20^{2} + 25^{2}\right) -2\left( 20 x 25 \right)\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

or would the 50 be cos(50) at the end

OpenStudy (jdoe0001):

\(\large \color{green}{a}=\sqrt{\color{blue}{b}^2+\color{red}{c}^2-2(\color{blue}{b}\times\color{red}{c})cos(\color{green}{A})} \)

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

It came out negative again. Agh

OpenStudy (anonymous):

What did I do wrong? I know thats not the answer

OpenStudy (jdoe0001):

make sure your calculator is in Degrees mode

OpenStudy (anonymous):

it is

OpenStudy (jdoe0001):

lemme recheck

OpenStudy (jdoe0001):

I get a positive number :)

OpenStudy (anonymous):

lemme try afgain

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

ha.

OpenStudy (jdoe0001):

5? ahemm, a bit of a small side, is a bit bigger

OpenStudy (anonymous):

oops forgot to add in cos

OpenStudy (anonymous):

it came out with -610

OpenStudy (anonymous):

ahhhhh

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

let's see

OpenStudy (anonymous):

what is wrong with me

OpenStudy (anonymous):

I typed it in exactly the same

OpenStudy (anonymous):

√ (20^2+25^2) -2 (20x25) cos(50)

OpenStudy (jdoe0001):

$$\large{ 20^2+25^2=1025\\ 2(20\times 25)=1000\\ cos(50^o)=0.64} $$

OpenStudy (jdoe0001):

so, what do you get?

OpenStudy (anonymous):

2023.64

OpenStudy (anonymous):

44.9

OpenStudy (anonymous):

45! YAY

OpenStudy (jdoe0001):

hehee, you may wan to recall the "order of operations" for numbers :)

OpenStudy (anonymous):

Oh yeah...ha. So not 45?

OpenStudy (jdoe0001):

if one leg is 20 and the other 25, another side of 45, would yield an angle bigger than 50 degrees

OpenStudy (jdoe0001):

I assume you know the "order of operations"?

OpenStudy (anonymous):

yeah pemdas

OpenStudy (jdoe0001):

1025-(1000*0.64)

OpenStudy (anonymous):

385?

OpenStudy (anonymous):

but now I need to square root it, correct? Or no

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

which would be 19.6

OpenStudy (jdoe0001):

yep

OpenStudy (anonymous):

YAY